doglover:

y=1/5x -1 in standard form

AngeI:

Hello sorry, I went afk for a bit. Do you know how to convert slope intercept form to standard form?

doglover:

yes i think so i just don't know how to start the equation

AngeI:

May you please say what you think you know so I can see how you do? I'll assist where needed c:

doglover:

well i think that i should first add one to each side, than divided both sides by 1/5 and then get x

doglover:

witch equals to 5+5y=x

AngeI:

So you would have to subtract 1/5x from both sides y=1/5x -1 -1/5x+y=-1 Then you would multiply everything by -5 to get rid of the fraction x-5y=5

doglover:

ok, thank you! , it makes since to me know

AngeI:

I'm glad O: